OpenStudy (anonymous):

determine weather the graphs of the equation are perpendicular 7x-9y=4 7y-9x=2

OpenStudy (mathteacher1729):

Step 1: Find their slopes. Step 2: Multiply their slopes together (slope 1 ) times (slope 2) Step 3: Is the product of the slopes (-1) ? YES = Perpendicular. NO = not perpendicular.

OpenStudy (saifoo.khan):

simple formula: m1 x m2 = -1

OpenStudy (anonymous):

so yes they are

OpenStudy (anonymous):

?

OpenStudy (saifoo.khan):

just a sec, brb i will giv u steps.

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

http://www.twiddla.com/563630

OpenStudy (saifoo.khan):

there?

OpenStudy (anonymous):

yeup

OpenStudy (mathteacher1729):

Saifoo -- the slopes are NOT 7 and 9. :( line one: 7x-9y=4 line two: 7y-9x=2 To obtain slope, you must put it into y = mx + b form. For line one: -9y = -7x + 4 y = (7/9)x + (4/-9) slope = 7/9 For line two 7y = 9x + 2 y = (9/7) + (2/7) slope = 9/7 So multiply (7/9) and (9/7) to obtain +1, which is NOT (-1), so the slopes are not perpendicular. See attached for pic.